State college has 620 students. There are 20 more women than men. How many women are there?
Found 2 solutions by Earlsdon, broske:
Answer by Earlsdon(6294)

If W = the number of women, then 620-W = the number men (M), and W+M = 620, and W = M+20 so...
W+M = 620 and W = M+20 or M = W-20, so...
W+(W-20) = 620
2W-20 = 620 Add 20 to both sides.
2W = 640 Divide both sides by 2.
W = 320 and M = 300
The number of women is 320.

Answer by broske(1)

total students - extra women / number of groups(M,W) = # of men
(620 - 20) / 2 = 300
number of men + extra women = total # of women
300 + 20 = 320